Using the data provided for a company in Jackson, what is the price-earnings ratio?

The following information pertains to a company in Jackson. Assume that all balance sheet amounts represent both average and ending balance figures and that all sales were on credit. Use this information to answer the questions that follow.

Assets

Cash and short-term investments

$ 30,000

Accounts receivable (net)

20,000

Inventory

15,000

Property, plant, and equipment

185,000

Total assets

$250,000

Liabilities and Stockholders’ Equity   

Current liabilities

$ 45,000

Long-term liabilities

70,000

Stockholders’ equity—Common

135,000

Total liabilities and stockholders’ equity

$250,000

Income Statement   

Sales

$ 85,000

Cost of goods sold

45,000

Gross margin

$ 40,000

Operating expenses

(15,000)

Interest expense

(5,000)

Net income

$ 20,000

Number of shares of common stock outstanding

6,000

Market price of common stock

$20

Total dividends paid

$9,000

Cash provided by operations

$30,000

Group of answer choices

2.5 times

6.0 times

8.0 times

4.0 times
Answer #1

EPS=(Net income/Number of shares of common stock outstanding)

=(20,000/6000)=$3.33(Approx)

P/E ratio=stock price/EPS

=20/3.33

=6 times